On June 1, 20--, a depreciable asset was acquired for $5,400. The asset has an estimated useful life of five years (60 months) and no salvage value.

Using the straight-line depreciation method, calculate the book value as of December 31, 20--. If necessary, round your answer to two decimal places.
Answer #1

Answer)

The straight line depreciation formula for an asset is as follows:

Annual Depreciation Expense Formula

Where:

Cost of the asset is the purchase price of the asset = 5400

Salvage value is the value of the asset at the end of its useful life = 0

Useful life of asset represents the number of periods/years in which the asset is expected to be used by the company = 5 years

So annual depreciation expense= (5400 - 0)/5 = 1080 $

Here, The total depreciation to be calculated from the month of june 1 to december 31 = 7 months

Hence,

Accumulated depreciation = Annual depreciation expence × 7/12

Hence, Accumulated depreciation is 1080 × 7/12 = 630 $

Book value of asset as of December 31, 20-- is:

= Cost of asset - Accumulated depreciation

= 5400 - 630

= 4770 $

Therefore, the book value as of dec 31st = 4770 $